OpenStudy (anonymous):

Simplify: square root of 3 times square root of 21 square root of 24 square root of 63 3 square root of 7 None of the above

OpenStudy (anonymous):

@JonnyVonny

OpenStudy (anonymous):

@wendylisette

OpenStudy (anonymous):

@Kristen17

OpenStudy (therealmeeeee):

u dont want me to help u

OpenStudy (anonymous):

i do lol

OpenStudy (therealmeeeee):

the answer is None Of The Above

OpenStudy (therealmeeeee):

D

OpenStudy (anonymous):

thanks @TheRealMeeeee

OpenStudy (therealmeeeee):

ill take a medal

OpenStudy (anonymous):

how do i give you a medal

OpenStudy (therealmeeeee):

click Best Respose

OpenStudy (jonnyvonny):

\[\sqrt{3}*\sqrt{21}=\sqrt{63}=\sqrt{9}*\sqrt{7}=3*\sqrt{7}=3\sqrt{7}\]

OpenStudy (jonnyvonny):

@TheRealMeeeee , how did you obtain your answer?

OpenStudy (therealmeeeee):

@JonnyVonny im right your wrong

OpenStudy (anonymous):

idk whos right either @JonnyVonny or @TheRealMeeeee

OpenStudy (therealmeeeee):

i will explain it

OpenStudy (anonymous):

ll give medal to who ever is right

OpenStudy (therealmeeeee):

I KNOW IM RIGHT @JonnyVonny got me yesterday 37% on my test yesterday i did it over and i got 100%

OpenStudy (therealmeeeee):

go look at his medal and fans and his testimony r all good go look at mine and u will see whos better

OpenStudy (jonnyvonny):

How about post your answer, put it in a calculator, then I shall do the same with my answer, and we shall compare it to the sqrt(3)*sqrt(21)

OpenStudy (anonymous):

yeah do that

OpenStudy (therealmeeeee):

i got 14 fans and 29 medals

OpenStudy (anonymous):

ok @TheRealMeeeee your getting 30 medals know

OpenStudy (therealmeeeee):

boom ill put my stuff up anyways

OpenStudy (anonymous):

good job ill give another medal if you help me in another problem

OpenStudy (therealmeeeee):

just look at this i got the same but i went on open study to check my work ill post up the so u can see it

OpenStudy (jonnyvonny):

The site must be wrong or something, because the answer I get from my calculator shows that I'm correct; we can get someone in here that know what to do and decide.

OpenStudy (dranz10):

https://answers.yahoo.com/question/index?qid=20110117161556AA9j9iW I think @JonnyVonny is right

OpenStudy (dranz10):

LOOK!!!

OpenStudy (dranz10):

@TheRealMeeeee